By what factor would your weight change if Earth's diameter were tripled and its mass were quintupled?
1
Expert's answer
2012-10-30T11:11:35-0400
W = your weight
G = gravitational constant
M = mass of Earth
m = mass of you
r = radius of Earth

W = GmM/r²

if the diameter of Earth were halved, so would be its radius
so your weight with radius halved and mass quintupled would be

W_new = Gm(5M)/(½r)² = 5GmM/(¼r²) = 20W_old